Find the maclaurin series expansion for y=(8+ex)1/2 in ascending powers of x upto and including the term in x3 hence estimate (8+e0.54)1/2 to three decimal places


1
Expert's answer
2021-07-27T16:20:58-0400
"f(x)=(8+e^x)^{1\/2}"

"f(0)=(8+e^0)^{1\/2}=3"

"f'(x)=\dfrac{1}{2}e^x(8+e^x)^{-1/2}"

"f'(0)=\dfrac{1}{6}"

"f''(x)=\dfrac{1}{2}e^x(8+e^x)^{-1/2}-\dfrac{1}{4}e^{2x}(8+e^x)^{-3/2}"

"f''(0)=\dfrac{1}{6}-\dfrac{1}{108}=\dfrac{17}{108}"

"f'''(x)=\dfrac{1}{2}e^x(8+e^x)^{-1/2}-\dfrac{1}{4}e^{2x}(8+e^x)^{-3/2}"

"-\dfrac{1}{2}e^{2x}(8+e^x)^{-3/2}+\dfrac{3}{8}e^{2x}(8+e^x)^{-5/2}"


"f'''(0)=\dfrac{1}{6}-\dfrac{1}{108}-\dfrac{1}{54}+\dfrac{1}{648}=\dfrac{91}{648}"

"f(x)=3+\dfrac{1}{6}x+\dfrac{17}{216}x^2+\dfrac{91}{3888}x^3+O(x^4)"

"(8+e^{0.54})^{1/2}\approx3+\dfrac{1}{6}(0.54)+\dfrac{17}{216}(0.54)^2"

"+\dfrac{91}{3888}(0.54)^3\approx3.117"